Tracker Meeting

February 19, 2004

 

Action Items:

-        Robert and Jim: look at the sidewall need date in the Tower-A schedule

-        Mike Menning: review the thermal-strap RTV decision with Jack and Jeff

-        Ben: send Jim and email about where to ship the left-over prepreg

-        Robert: update the bias circuit NCR

-        Jerry: update the grounding tube in the tray assembly drawings

-        Sandro: get a time estimate from G&A on when they will have results from wirebonding the preproduction MCMs

-        Dave: ship 10 preproduction MCMs to Pisa

-        Jeff: prepare and ship 2 burn-in cables for Pisa

-        Jeff: talk to Gunther about how to ship items to Italy

-        Erik: complete the static test procedure document by next week

-        Sandro: find a person to be responsible for static testing

 

Agenda:

††††††††††† - Sandro & Erik: feedback on the redlined titanium parts drawings

††††††††††† - Jeff: information needed on sidewall thermal test coupons

††††††††††† - Nanda/Ben: Sidewall test matrix

††††††††††† - Mike: status of titanium parts drawings and procurement

††††††††††† - Dave: MCM status

††††††††††† - Jerry: bias circuit status (scrapping old circuits)

††††††††††† - Sandro: G&A progress/status on MCM integration

††††††††††† - Riccardo: Plyform status on mid tray production and bottom-tray assembly tooling

††††††††††† - Nicola or Jack: T/V test status

 

Last week's action items:†††††††

††††††††††† - Nanda: Status of prepreg procurement

††††††††††† - Jeff: Mid-tray assembly drawing status (grounding tubes)

††††††††††† - Sandro: ownership of static testing

††††††††††† - Tom V: bakeout specs

 

Nicola reported that work is in progress assembling the tower with thermo-couples.They plan to release procedures and plans tomorrow for the Feb 26 TRR.Sandro reported that the resistance seen by each MCM is now well known.They will put the TEM on the tower when the tower is still open and will power it up and probe at different points the voltage.Sandro also reported that they have all the pieces for the blanket structure and Al box.

 

Jeff and Jack need information on the thermal coupons.Each group of 6 coupons, now at SLAC, is identified by a part number and group number, but SLAC does not know what those numbers mean.Riccardo has the information but is at Plyform.He should have the IDs of the coupons.There should also be a drawing from Tom Borden.

 

Nanda reported that Plyform reviewed the prepreg specs and test matrix.They had questions on the test matrix.Specifically, is it necessary to test coupons from every sidewall?This would require a lot of extra material.Plyform would like to do test matrix once per purchase lot and keep coupons for all sidewalls from the border.Full thickness test coupons come from a separate layup.

Ben's proposal: do the full test matrix for every material purchase lot.Do in addition the short-beam shear test on every cure.Those coupons are small, 20mm by 10mm, and can be cut off close to the sidewall.Also, the fiber volume and void content need to be measured on each layup. ††Each sidewall will also get a dimensional inspection.

 

Nanda said that Plyform is moving forward with the spec toward purchasing prepreg.Robert and Jim need to take a look at the schedule constraints.Sandro estimated for the manufacturing time 5 or 6 weeks total, after the prepreg is received.It was suggested that INFN should use SLAC to expedite response from COI if necessary.

 

Erik said that he looked over the redlined drawings from Mike and Martinís meetings in Pisa and went over the analysis with John Ku.Erik had discussed some concerns with Mike Menning.Mike Foss will have the drawings done this afternoon.

 

Mike made some comments to address Erik's concerns on the flexure blade fit into the corner bracket.It is a transitional press fit in both dimensions.He will send the drawings to the vendors, then travel there and discuss details of tolerances.If a light press fit is thought by the vendors to be possible, then they will negotiate that.

 

Mike sent Ron in purchasing an email asking for a meeting tomorrow morning to adjust details of the titanium PO.He will add source inspection of the first two articles before authorizing the remaining items.Jeff will help Mike make sure there are no administrative hangups.

 

Tom Venator reported that the Goddard expert (Chris) said that from a contamination standpoint there is no requirement for a tray bakeout.Tom will talk to him more in detail, but he suggested that if the bakeout is a workmanship test, then 24 hours is not necessary.Sandro asked if there were bakeout requirements regarding the Hysol EA934 glue cure. ††Ben says it is commonly used with a 7-day room temperature cure, so no bakeout is required for that reason.We will go ahead with the bakeout as a pre-silicon workmanship test, but cutting the time down could help our schedule.

 

Mike said that Riccardo owes a sketch of the shipping container and lifting fixture and also a redlined drawing for the thermal strap.Sandro replied that Emelio has prepared sketches to send tomorrow for the container and fixture.Nothing has been done yet on thermal straps.

 

Jeff asked if there were Italian governmental regulations on the lifting fixture.None were known, but we decided that it would be safe to design it to NASA requirements.

 

Mike asked about use of RTV on the thermal strap where it gets sandwiched between Tracker sidewall and bottom tray.This is not presently in the drawings.We discussed what material to use and settled on the electrically conductive Nusil that is already being used to glue down ladders and to glue the grounding tube into trays.

Mike will review this decision with Jack and Jeff.

 

Jim asked about disposition of COI excess material from the sidewall build, which is about to expire unless we pay to requalify it.SLAC has decided to scrap it.There are 22 square feet of YS90 and several pounds of K13D.Ben will send Jim email about where it could be sent at GSFC.

 

Sandro reminded us that 4 completed sidewalls are already at Plyform (Robert pointed out that another 4 are at COI).We will keep in mind the option to use these by post-bonding 25micron aluminum to the inside face and redrilling the countersinks.Sandro estimates a month lead time for that.He will have to ask Plyform about the post-bonding operation, whether we can expect any problems with that.

 

Jerry reported on the bias circuits.He needs to get modifed Gerber file from Tung (this was changed to Thien later in the day).Robert and Sandro will review the resulting Gerber data.A 3D model is being made and will be checked against the import Gerber files.We will scrap all of the existing circuits.†† New circuits will have the serial number, will not have extended traces for plating, will have the circuit properly centered w.r.t. the alignment holes, and will have grounding pads modified to give much better clearance from the high-voltage SSDs.Darren pointed out that the NCR needs to be completed.

 

Jerry reported that he has taken the action for marking up the tray assembly drawings for the grounding tubes and getting the drawings re-finalized.

 

Sandro reported that G&A has the tray mockup drawings and 10 MCMs.They agreed with him on a scheme of mechanical measurements.Sandro will ask them for a time estimate to complete this work.He had discussions with them for a tool to attach MCMs to trays. Next week he will send a sketch.The tool includes a good idea how to recover alignment of the pitch-adapter to the bias circuit and SSDs.

 

Dave reported on the MCM status.

Preproduction: 14 were completed a while ago (10 of which are in Italy).Another 14 recently completed burn-in and in final test and others are in progress at Teledyne. Teledyne started Tuesday the final 10 with V7 chips, using tall boards.We expect 2 weeks to complete this build cycle.

Flight production: SLAC screened 150 pitch adapters so far and rejected about 12%.Dave is adding temperature and humidity monitoring, because we observed systematic differences in average length from one week to the next.Dave and Jeff are kitting flight materials at SLAC.Tomorrow they will ship about 25% of the flight build for all parts available.†† The gating item is coupon-tested, machined PWBs.This has been held up partly by a need to qualify a new machine shop, as Holt will no longer do this work for us.25 are at a new vendor today.They will get CMM at SLAC when done.Our plan is to start flight production next week.We still need to close on the PO, and there is a meeting for that tomorrow.Teledyne on their side has been reprogramming the wire bonding so that all wires bond to one side of the ASIC pad, leaving undisturbed metal for making a rework bond when necessary.This will allow all rework wire bonding to be done to Mil Spec.

†††††††††††

Sandro requested that SLAC send 10 preproduction MCMs and 2 burn-in cables to Pisa, for software development.Jeff said there is a problem with getting connectors to assemble more burn-in cables right now, but Robert and Jeff will try to find enough parts for 2 cables.

 

Sandro reported a shipping problem.VME processors arrived in Italy FedEx.This causes a big delay and hassle in customs.Jeff Tice will talk to Gunther about this and request that all shipments to Italy go through him.Jeff knows how to ship items to the Pisa airport, cargo-no-courrier.

 

The static test system left for Italy last week. Erik will have his procedure document done in 1 week.Sandro will hold a meeting to decide on the responsible person in Italy for this equipment and test.After these two items are taken care of we can consider the need to send Erik to Italy for setup and training.

 

Honeycomb is being inspected today and tomorrow in Plyform.Plyform is generating a document which will be appended to the NCR.

 

Sidewall thermal cycles were done at Terni and the results were good, with no effects seen on the walls.This has been documented in LAT-TD-03065.